Trying to develop a linear equation for:
Gas used: 5, 8, 12, 15, 18, 20, 22, 25
Miles driven, 80, 100, 180, 225, 280, 290, 320, 375

Answer by drk(1908)

We have the following table:
Gas used: 5, 8, 12, 15, 18, 20, 22, 25
Miles driven, 80, 100, 180, 225, 280, 290, 320, 375.
To create a linear model, we want to apply some rule to x and then get y.
Notice
5*15 = 75 + 5 = 80
8*15 = 120 - 20 = 100
12*15 = 180
15*15 = 225
18*15 = 270 + 10 = 280
20*15 = 300 - 10 = 290
22*15 = 330 - 10 = 320
25*15 = 375.
We can't create a perfect linear model, but this means we have a "regression" model, or an equation that gets us close:
M = 15G.
Hope this helps.